An engine raises 100kg of water through a height of 60m in 20s. What is the power of the engine?
(g= 10m/s)
1
Expert's answer
2019-07-04T09:24:20-0400

We can find the power of the engine from the formula:


"P = \dfrac{W}{t} = \dfrac{mgh}{t} = \dfrac{100kg \cdot 10 \dfrac{m}{s^2} \cdot 60 m}{20s} = 3.0 \cdot 10^3 W = 3.0 kW."

Answer:

"P = 3.0 \cdot 10^3 W = 3.0 kW."